### v1

# Mapping atmospheric features of the planetary-mass brown dwarf SIMP 0136 with JWST NIRISS

Abstract
In this paper, we analyze James Webb Space Telescope Near Infrared Imager and Slitless Spectrograph time-series spectroscopy data to characterize the atmosphere of the planetary-mass brown dwarf SIMP J01365662+093347. Principal component analysis reveals that 81% of spectral variations can be described by two components, implying that variability within a single rotational phase is induced by at least three distinct spectral regions. By comparing our data to a grid of Sonora Diamondback atmospheric models, we confirm that the time-averaged spectrum cannot be explained by a single model but require a linear combination of at least three regions. Projecting these models onto the principal component plane shows that the overall variability is highly correlated with changes in temperature, cloud coverage, and possibly effective metallicity. We also extract brightness maps from the lightcurve and establish North-South asymmetry in the atmosphere. A combined multidimensional analysis of spectro-photometric variability links the three spectral regions to three atmospheric layers. Forsterite cloud and water abundance at each level form unique harmonics of atmospheric variability observed in different spectral bands. Atmospheric retrievals on the time-averaged spectrum are consistent with an optically thick iron cloud deck beneath a patchy forsterite cloud layer and with the overall adiabatic curve. We also demonstrate two new analysis methods: a regionally-resolved spectra retrieval that relies on multi-wavelength spherical harmonics maps, and a method to constrain brightness maps using Doppler information present in the spectra. Future observations of variable brown dwarfs of higher spectral resolution or spanning multiple rotations should help break mapping degeneracy.
show]roman.akhmetshyn@mail.mcgill.ca
1 Introduction
Rotational variability is common among brown dwarfs (Artigau et al., 2009; Radigan et al., 2012; Buenzli et al., 2014; Radigan et al., 2014; Metchev et al., 2015; Lew et al., 2016; Artigau, 2018), planetary mass objects (Biller et al., 2015, 2018) and directly imaged companions (Zhou et al., 2016; Naud et al., 2017; Zhou et al., 2020a, b), especially those of young age and low surface gravity (Vos et al., 2018, 2023). Changes in lightcurve morphology are commonly attributed to the dynamic formation and dissipation of clouds, radiative localized hot spots, intense vertical transport due to chemical instability, or the differential rotation of clouds at different latitudes (Bailer-Jones & Mundt, 2001). Specific morphology during a single rotation may be due to cloud features rotating in and out of view. Outstanding questions include whether the atmospheric structure is dominated by banding, spots, or a mixture of both; and measuring the timescales of atmospheric variability to constrain the driving mechanisms.
SIMP J01365662+0933473 (SIMP 0136 hereafter) is a nearby ( pc; Gaia Collaboration et al. 2023) T2.5 brown dwarf discovered in 2006 at the Observatoire du Mont Megantic with the wide-field near-infrared camera CPAPIR as part of SIMP (Sondage Infrarouge de Mouvement Propre) proper-motion survey (Artigau et al., 2006). Further observations revealed a rotation period of hr, a peak-to-peak -band amplitude of 50 mmag, and significant evolution of lightcurve between each observations (Artigau et al., 2009; Croll et al., 2016; Yang et al., 2016; Plummer et al., 2024). These properties make SIMP 0136 the brightest isolated T dwarf and an ideal target to study brown dwarf atmospheric variability. Measurements of constrained its inclination to ∘ (Vos et al., 2017). Further kinematic measurements identified SIMP 0136 to be part of 200 Myr Carina-Near moving group which allowed to estimate its mass of MJ using brown dwarf evolutionary models (Gagné et al., 2017). The same work also estimated that SIMP 0136 could have already depleted half of its deuterium content. Given its mass at the planet-star boundary, SIMP 0136 has been called a free-floating planet. Multiple ground- and space-based observations confirmed its variability across multiple rotations (Artigau et al., 2009; Croll et al., 2016; Yang et al., 2016; Plummer et al., 2024; McCarthy et al., 2025). General circulation models (GCMs) predict that this variability is mainly caused by patchy cloud layers (Tan & Showman, 2021a). It is believed that different cloud decks form and dissipate at different time scales, rotate in and out of the view and hover over each other in different configurations (Apai et al., 2013; Metchev et al., 2015; Apai et al., 2017; Vos et al., 2023).
Section 2 describes the data and its reduction. In Section 3 we show the principal component analysis of our data. In Section 4 we compare NIRISS time-series spectra to atmospheric models. In Section 5 we describe our atmospheric retrieval and compare with the results from previous works. Section 6 describes different mapping approaches, focusing on the results from spherical harmonic mapping. Section 7 explains our multidimensional study of spectral variability and connects lightcurve morphology with meteorological structure. Section 8 explains how spherical harmonics maps can be used to perform extraction and retrieval of regional spectra. Section 9 proposes to constrain brightness maps with Doppler measurements using JWST spectroscopy.
2 Time variable data
The data for this research were obtained with the Near Infrared Imager and Slitless Spectrograph in the Single Object Slitless Spectroscopy mode (NIRISS/SOSS) onboard the James Webb Space Telescope’s (JWST) as a part of cycle 1 Guaranteed Time Observations program #1209 (PI: Artigau). Observations were carried out from UT 05:08:28 to 09:05:07 on July 22, 2023, 37 hours before JWST Cycle 2 program GO #3548 (PI: Vos) that observed SIMP 0136 in NIRSPec/BOTS (McCarthy et al., 2025). We used the SOSS-Inspired SpectroScopic Extraction (SOSSISSE) pipeline for data reduction, which was previously used and described in Lim et al. (2023).
The reduced data consists of 81 1200 spectra from 0.85 to 2.83 m taken over slightly more than one rotational period (Figure 1). The spectrum changes during the object’s rotation, and this variability is different for each wavelength bin associated with certain molecules, suggesting different meteorological structure at different depths.
3 Principal component analysis of time-varying spectra
We use principal component analysis (PCA) to find orthogonal axes that capture the most variance in the spectrum of SIMP 0136 . Since we have 81 spectra, 81 principal components can perfectly replicate all the information in the data. We found that binning data in time does not increase the fraction of variance captured by the first 2–3 of principal components. On the other hand, binning in wavelength greatly increases the fraction of variance explained by the first 2 principal components. Unbinned data have a lot of uncorrelated noise distributed over many eigenvectors, presumably Poisson noise. Averaging it out helps to focus on astrophysical variability. In the rest of this section we use 0.019 m binning, which reduces the number of wavelengths to 93, while the number of principal components remains at 81.
About 81% of the spectral variability can be explained by just two principal components. This justifies projecting our data onto the principal component plane (PCP) made of the first two PCs. According to Cowan et al. (2011), the number of dimensions explaining brightness variability determines the minimum number of distinct spectral regions contributing to the variations: . These spectral regions could represent different cloud decks or clear patches.
The PCP is a useful tool for interpreting the time evolution of spectra (Fuhrmann, 1999; Cowan et al., 2013). In Figure 2, each datum point represents a full spectrum at a given time. From Cowan & Strait (2013), if different spectral regions combine linearly to determine the planet’s overall spectrum, then the pure spectrum of each region must define vertices of a simplex that encloses the data. A simplex is a polyhedron with vertices. For our 2-dimensional PCP a simplex is a triangle. However, as seen in Figure 2, the projected spectra hardly represent a triangular structure and would be better enclosed with a quadrilateral. This suggests that we have 4 end-member spectra.
4 Comparison with Sonora Diamondback Atmospheric Models
In this section we follow a similar approach to Vos et al. (2023) and McCarthy et al. (2024), exploring different possible combinations and number of models that best fit our NIRISS observations. We utilize Sonora family Diamondback spectra models — a new grid of atmospheric models for substellar objects with a range of parameters like temperature, cloud coverage and metallicity (Morley et al., 2024a). We use a limited temperature grid range based on our prior knowledge of SIMP 0136. We also do not constrain gravitational acceleration, because previous studies report values between 100 and 316 cm s-2, see Table 1 for grid parameter values that we explored.
Before performing model matching, we compare the integrated flux of Diamondback models to , where is the reported temperature of a given model spectrum. We find that the large majority of grid spectra have higher integrated temperatures than reported for each model. For our grid range, the average difference is 25 K, the maximum difference for 900 K objects is 3%, which rises to 5% for 1400 K objects. We do not find these discrepancies to be an issue, as they are comparable to derived temperature uncertainties from other studies. However, one must be careful when using grid models to infer the target’s temperature.
| T (K) | g (cm s-2) | [M/H] | |
| 900 | 31 | 1 | |
| 1000 | 100 | 2 | 0.0 |
| 1100 | 316 | 3 | |
| 1200 | 1000 | 4 | – |
| 1300 | 3160 | 8 | – |
| 1400 | – | NC | – |
Note. is the sedimentation efficiency parameter that characterizes cloudiness: smaller number means more clouds. NC refers to clear-atmosphere spectra. All models use [C/O] = 1.0.
References. Morley et al. (2024b).
Rescaling the observed spectrum to the top of the atmosphere flux requires knowledge of the radius of the object and the distance to it. While distance is precisely known from Gaia EDR3 (Gaia Collaboration et al., 2023), radius estimations for brown dwarfs are usually inconsistent (Burrows et al., 2011). Atmospheric modeling with BT-Settl (Allard et al., 2013) predict SIMP 0136 to be 0.82 RJup, while evolutionary model SM08 (Saumon & Marley, 2008) given the age estimation provides 1.22 RJup (Gagné et al., 2017). In all our calculations, we adopt the latter, as it gives the most comparable fluxes to Diamondback spectra for the most probable T range. A decrease in radius would greatly increase the emitted flux, thus increasing spectrum-inferred brightness temperature and T of the best-fit Diamondback model above widely accepted values.
First, we search for a single model that best fits the time-averaged spectrum. This corresponds to a model with T K, gravitational acceleration g cm s-2, f, and metallicity of 0.5. To improve the accuracy, we then create finer-grid interpolations around the best model. The best interpolated spectrum has the following parameters: T K, g cm s-2, f, and metallicity of 0.2. As seen in Figure 3, the interpolated model poorly fits the region around 2.0 – 2.5 m, the 1.25 m and 1.65 m peak, and 1.4 m water absorption band. None of the other single grid models described well the time-average spectrum or spectrum at any timestamp. This suggests that the observed spectra are linear combinations of multiple distinct spectra.
We used the Bayesian Information Criterion (BIC) to analyze how many models are needed to best describe our data without over-fitting. The fitting was performed iteratively: starting with the best-fitting single interpolated model, we added additional models and recalculated weights one at a time to improve the overall fit. Our The NIRISS spectrum is incredibly precise, and to adequately apply the BIC, we inflated the uncertainties to match the standard deviation of the spectral variability over time. The number of parameters is , and the BIC is minimized for . This could mean that the observed time-average spectrum is an admixture of spots that have different temperatures and chemistry. An atmosphere explained by 6 spectral models is in agreement with previous PCA results, where majority of the variance is explained by 2 components, and thus there are spectral regions that change in time. Combined PCA and BIC analyses are illustrated in Appendix A. However, without inflating the observational uncertainties, no linear combination of models provides an acceptable match to the data, suggesting that the current models are missing some physical processes and highlighting the need for further refinement.
We constrain the number of admixture components to three based on the results of the PCA and find a best-fit combination for matching the time-averaged spectrum. Physical parameters of these 3 models are: 1000, 1200, and 1300 K; 3, 2, and 8 fsed, 100, 100, and 316 cm s-2; -0.5, +0.5, and +0.5 [M/H]. The integrated Stefan-Boltzmann temperature of the admixture is 1168 K. As shown by the red curve in Figure 3, this reduced admixture model performs only marginally worse than the more complex six-component fit to the time-averaged spectrum. We then use this 3-model combination to fit the time-varying spectra. In Figure 4, we assess whether the observed spectroscopic variability can be captured by the three-model fit. While it slightly underestimates the flux at certain phases, the overall morphology of the variability is well reproduced.
5 Atmospheric retrievals and T-P profiles
We performed atmospheric retrievals on the time-averaged spectrum using the Brewster framework (Burningham et al., 2017, 2021). The temperature profile was parameterized following Madhusudhan & Seager (2009). Our model accounts for the opacity contributions of H2O, CO, CO2, CH4, NH3, CrH, FeH, Na, K, which are spectroscopically active in the near-infrared and have been previously identified as key absorbing species in the spectra of L/T transition objects. Additionally, continuum opacities from collision-induced absorption and Rayleigh scattering due to H2, He, and CH4 are included. Previous retrieval analyses of SIMP 0136 (Vos et al., 2023) indicate the archival 1–15 spectral data are best described by a patchy, high-altitude forsterite/enstatite (Mg2SiO4/MgSiO3) cloud overlaying a deeper, optically thick iron (Fe) cloud; we follow this prescription.
Figure 5 presents the retrieved temperature profile and cloud distributions, compared to Diamondback models (the same used in Figure 3) and phase-equilibrium condensation curves. The retrieved temperature profile (brown curve) shows an adiabatic slope below 1 bar, in good agreement with retrieval results of Vos et al. (2023) that combine 1–2.5 m Infrared Telescope Facility /SpeX Prism spectrum, a 2.5–5 m AKARI/Infrared Camera spectrum, and a 5–15 m Spitzer/IRS spectrum. Self-consistent grid models of the best-fit 3-spectra admixture model are also plotted for reference with Teff = 1000, 1200, 1300 K, g = 100, 100, 316, and = 3, 2, 8. While being broadly consistent, our retrieved profile is more isothermal at high altitudes and slightly cooler at 1–10 bars. This discrepancy may be attributed to the limited flux contribution from the upper atmosphere in the near-infrared bands, resulting in weaker constraints on the temperature structure at these altitudes.
6 Mapping
We may infer some brightness geometry of stars and planes via observed light curve variations. This inference is complicated by the fact that the brightness-to-map problem is degenerate: one brightness map results in a single light curve, but one light curve can be produced by potentially infinite map variants. Knowledge of rotational period, inclination and limb darkening can help to constrain some parameters.
In this work we use two complementary techniques to map SIMP 0136: sinusoidal (Cowan & Agol, 2008; Cowan et al., 2013), and spherical harmonic mapping (Cowan et al., 2013; Luger et al., 2019). The first method is algebraically simple and assumes a north-south symmetry of the object. The second uses spherical harmonic basis maps.
If the thermal map is North-South symmetric, then odd harmonics do not have a phase function signature. Instead, our observations of SIMP 0136 exhibit significant power at the 3rd mode (see Figure 6b), indicating North-South asymmetry and inclined rotation (Cowan et al., 2013).
Spherical harmonics within the Starry framework (Luger et al., 2019) allow to account for the inclination of the object. We provide a multivariate Gaussian prior on the spherical harmonics coefficients with a mean and covariance.
The results from fitting the m band lightcurve with various mapping methods are shown in Fig. 6a. A sinusoidal fit with 4 modes, excluding the 3rd mode, is represented by the dashed green line. The maximum number of modes was selected using Bayesian Information Criterion (BIC). This approach failed to capture the distinct shape of the phase curve, nor similar features in other bands. If we include the odd harmonics (dashed blue line), we get a reasonable fit. This means that SIMP 0136 has a North-South asymmetric map.
Spherical harmonic decomposition accounts for inclination and produces 2D maps. Using BIC we constrained number of -degrees to 4, that is 25 Ylm harmonics in total. We provided 2 sets of priors on coefficients: flat with uniform variance, and a spot-band structure. The resulting maps can be seen in Fig. 6c.1 and c.2. The asymmetric checker-board pattern is the most favourable for flat priors. A more plausible map is obtained with a spot-band prior. The fact that these maps produce identical lightcurves is a testament to the extreme mapping degeneracy (Cowan et al., 2013).
7 Connecting vertical structure to lightcurve morphology
SIMP 0136’s spectrally inhomogeneous variability indicates multiple layers with distinct atmospheric processes and dynamics. The enhanced NIR spectrophotometric variability observed at the L/T spectral transition has been demonstrated to be best fit by Fourier modes for objects such as WISE 1049B (aka Luhman 16B), VHS 1256 b, and SIMP 0136 (e.g., Apai et al., 2017, 2021; Fuda et al., 2024; Zhou et al., 2020a; Plummer et al., 2024). In this section, we employ a contribution function (and our previous harmonic analysis above) to map spectral wavelength ranges to pressure levels and create 2D (pressure vs time) flux variability maps which can be decomposed to determine variability sources at each atmospheric layer.
First, we fit the multi-spectral lightcurves with Fourier models using Imber (Plummer & Wang, 2022, 2023) which employs dynamic nested sampling (Skilling, 2004, 2006) via the open source Python code Dynesty (Speagle, 2020). Considering the spectral band from , we use bins to plot 36 light curves (see Figure 1 for an example with 0.2 bins). Qualitatively, three characteristic light curve shapes can be seen in these results, broadly matching the findings of Biller et al. (2024) and Chen et al. (2025) for WISE 1049B and McCarthy et al. (2025) for SIMP 0136 which found 3, 4 and 9 characteristic light curve shapes, respectively, with JWST/NIRSpec data. For each binned light curve, we select the best-fit Fourier model based on the Bayesian Information Criterion.
To create vertical maps, we use the best-fit Fourier models and a 1150 K, (Vos et al., 2023; Gagné et al., 2017), cloud-free Sonora Bobcat substellar atmospheric model (Marley et al., 2021) and Picaso (Batalha et al., 2019). At each time-step, we scale 0.05 m interval spectral bins in the contribution function by the corresponding normalized light curves from Figure 1. We then sum the total emission flux at each pressure level across all wavelengths. This creates a 1D vector of fluxes as a function of pressure. We repeat this process at every time step and then normalize the flux at each pressure level. The resulting vertical flux variability maps (see Figure 7) suggest three distinct layers, which match the qualitative light-curve morphology groups seen in Figure 1.
To analyze each layer’s unique harmonic components, we consider the specific contribution from each of its three layers by wavelength (, , and ). We do this for each layer by subtracting out the contributions from the other two layers. Each layer’s unique atmospheric flux variability map can be seen in Figure 7a, b, and c.
The deepest layer () corresponds to (Figure 7a), includes higher-order harmonics (, where is the wave number), and corresponds to the spectral region where the models predict cloud variability (e.g., McCarthy et al., 2025). We interpret higher-order harmonics at this pressure level as forsterite (Mg2SiO4) cloud modulation, in line with Vos et al. (2023), McCarthy et al. (2024), and Plummer et al. (2024).
The transition layer () at 1 bar corresponds to the expected pressure level of Mg2SiO4 cloud tops (e.g., Vos et al., 2023; McCarthy et al., 2024) and exhibits a transition from waves including complex, higher-order harmonic waves to those dominated by lower-order harmonics. The layer contains bright spots correlated with the suspected cloud features at the top of the deeper layer. This layer corresponds to a prominent water absorption band and we suspect the bright spot in Figure 7b depicts either localized heating or variations in observed water absorption tied to the presence of forsterite clouds, either due to muting of spectral features or the depletion of water as part of the thermochemical reaction that produces forsterite clouds (Visscher et al., 2010).
The highest altitude layer () with pressures mbar (see Figure 7c) typically contains lower-order harmonics (). This high layer’s wavelength range includes both water and carbon monoxide (CO) absorption bands. As can be seen by comparison to Figure 7c, the upper layer’s flux variability is in phase with the bright spot in the transition layer at bar.
8 Regional spectra retrieval
Cubillos et al. (2021) proposed to invert phase-dependent spectral observations and produce longitudinally resolved spectra that can then be fit using standard 1D spectral retrieval codes. Since SIMP 0136 ’s variability is more accurately described by spherical harmonics maps, we extend the method of regionally resolved spectra to 2D maps. By having time-resolved spectra covering the full rotational phase, it is in principle possible to construct brightness maps at different wavelengths and extract spectra from specific regions on the map.
In our case, the extracted regional spectra and their corresponding atmospheric retrievals are unreliable due to degeneracies inherent in 2D mapping. We also encountered difficulties due the absence of grid-based flux in the forward model. As already shown by Cubillos et al. (2021), the grid flux normalization is not trivial. Nevertheless, some results can be interpreted. The temperature structure varies significantly across different regions, suggesting that variability is primarily due to temperature inhomogeneity. Some species are well-constrained and consistent between regions: water, methane, alkalies, iron hydride. However, regional spectra show higher water, alkali, and methane abundances than time-averaged spectrum. At the same time, CO and carbon oxide, ammonia, chromium hydride have very different and uncertain posterior distributions. This could suggest that carbon chemistry is not the main source of variability. Additional details on the regional spectral retrieval attempts are provided in Appendix B.
Although this approach did not yield robust results, we consider it valuable to document the steps taken and challenges faced, as they may inform future efforts in this area. Notably, for objects exhibiting North–South atmospheric asymmetry, spherical harmonics maps may remain the only viable method to study individual spectral regions—provided that sufficient observational constraints are available.
9 Doppler Tomography
While the spectral resolution of NIRISS/SOSS is modest compared to ground-based precision radial velocity (pRV) spectrographs, the instrument’s stability enables meaningful Doppler measurements. The SOSSISSE framework describes the data as fractional variations relative to a mean flux, presented as a residual map. To measure velocity shifts, we adopt the approach of Bouchy et al. (2001), in which radial velocity shifts are determined by projecting the residuals onto the derivative of a high-SNR template. Although originally developed for the pRV field, this method is applicable to any spectroscopic time series. The Bouchy et al. (2001) formalism also provides associated uncertainties (see their Eq. 13).
Given the low spectral resolution of SOSS and its limited constraints on radial velocities, we do not attempt to divide the domain into wavelength bins as is done with flux variability. Instead, we consider the integral of Eq. 9 from Bouchy et al. (2001) across the 1.00–2.35 m range. The redder portion of the spectrum is excluded due to its markedly different light curve behavior.
Figure 8 shows the resulting radial velocity time series, revealing correlated variations at the km/s level, characterized by a three-harmonic signal. Per-point uncertainties are approximately 1 km s-1. The signal is in phase with the time derivative of the flux, consistent with the framework described by Aigrain et al. (2012), in which RV variability for a rotating object (neglecting convective inhibition) scales with the product of the flux and its time derivative—the so-called term in activity-related RV corrections.
The amplitude of RV variations observed for SIMP0136 is consistent with expectations based on its basic physical properties. For a rotating sphere with surface inhomogeneities (e.g., spots), one expects RV jitter on the order of the projected rotational velocity () multiplied by the relative photometric variability. With percent-level variability and km s-1 (Vos et al., 2017), the observed RV jitter matches the expected order of magnitude.
These results underscore an unexplored opportunity in JWST time-series spectroscopy of brown dwarfs. The high precision of molecular-band flux variability measurements—informing cloud structures at different altitudes—can be complemented by RV-based constraints on longitudinal and latitudinal features. While phase-resolved light curves constrain longitudinal structures, RV amplitudes are particularly sensitive to low-latitude features, providing strong priors for brightness map reconstructions. Although the SOSS data for SIMP 0136 lacks sufficient SNR to allow subdivision by bandpass, higher-resolution modes such as NIRSpec’s G140H or G235H are of particular interest. For a profile limited by instrumental resolution, RV uncertainty scales with the 3/2 power of resolution at a fixed flux level. Thus, moving from (SOSS) to (G140H/G235H) yields a theoretical sevenfold improvement in SNR, enabling subdivision into several tens of spectral segments while retaining km/s-level RV precision.
10 discussion and conclusion
In this work, we presented a multifaceted analysis of time-resolved spectroscopy of the planetary mass object SIMP J01365662+093347. The data were acquired with JWST NIRISS/SOSS and reduced with SOSSISSE. As expected from JWST’s stability, and amply demonstrated with transit work, the data is far beyond what could be obtained through ground-based observations of BDs. This level of detail opens a new window into BD atmospheric dynamics.
PCA showed that there are at least 3 different spectral regions rotating in and out of view. Indeed, not a single atmospheric model could fit observed spectrum, and a mixture of 6 Diamondback models is needed to adequately fit the time-averaged spectrum. Projected interpolated Diamondback models on the PCP suggest that disk-integrated physical parameters vary in time. These parameters are temperature, metallicity, and sedimentation efficiency. The fact that we have small spectral variations and change in integrated physical parameters in time ( K, f, ), suggests that SIMP 0136 has small-scale cloud patches of different temperature, as predicted by brown dwarf GCM results (Tan & Showman, 2021b). Discrepancy in the reported temperatures of Diamondback models does not affect the magnitude of temperature variability as it only creates a linear bias. The inhomogeneity in the integrated effective metallicity may arise from different amounts of chemical disequilibrium in each spectral region, e.g., differences in the CO/CH4 ratios among cloud layers.
Atmospheric retrievals of NIRISS spectra produce T-P profiles that deviate from ground-based and Spitzer retrievals in Vos et al. (2023). We suspect this is due to a more limited wavelength coverage of our dataset. Results of both retrievals confirm an iron cloud deck deeper in the atmosphere nevertheless.
Odd harmonics in the lightcurve of SIMP 0136 and the results from spherical harmonics mapping imply North-South asymmetry, combined with the previously noted inclined rotation. The 2D maps inferred from a single rotation remain degenerate. Nevertheless, lightcurve morphology in combination with Sonora Bobcat contribution functions reveal multidimensional structure of observed variability. We find that the spatial distribution of forsterite clouds below the 1-bar level is anti-correlated with the abundances of water and CO at higher altitudes. This is evidenced by a decrease in disk-integrated emissivity from forsterite layers, accompanied by an increase in emissivity from layers associated with water and CO absorption bands. North-South asymmetry of the brightness map in the domain could directly arise from the spatial distribution of small-scale forsterite cloud patches. It is not guaranteed that this weather pattern is a permanent feature on SIMP 0136: multi-period back to back observations could show how this morphology evolves.
The fact that the meteorology of SIMP 0136 evolves is confirmed by several ground- and space-based observations at different epochs. In Artigau et al. (2009), -band (m) photometric observations at Observatoire du Mont-Mégantic in 2008 reveal how the power of Fourier modes varies across nights. Phase-folded lightcurves from September 18th and 19th can be well explained by the first 2 modes, while phase curves of September 16th and 21st require an odd 3rd mode harmonic. Further -band observations at Lowell Observatory in 2016 (Croll et al., 2016) show consistent single-peak lightcurves across 4 rotations on the night of November 10th. On the night of November 12th, SIMP 0136 shows almost no variability in -band, indicating that meteorological features have longitudinally homogenized. The next night, however, sawtooth-like variations are consistently observed again, but in an anti-correlation with the night of November 10th, which suggests a slightly denser distribution of clouds in one hemisphere. Plummer et al. (2024) analyzed CFHT -band observations obtained in 2012: the lightcurve from October 14th is described by 3 Fourier modes; however, the following nights observations alternated between Y, J, H, and K bands, limiting the cadence and only allowing 2 modes to be inferred for each band. A phase shift between H-K and the J-H and H-K color indices was found, which might have indicated different cloud layers. Lastly, we compare our observations with JWST GO #3548 NIRSpec/BOTS time-series spectroscopy analyzed by McCarthy et al. (2025). The binned lightcurve in 0.9 – 2.1 m range is well explained by 3 Fourier modes, with the 3rd mode having equal power as the 2rd mode, just like in our observations. Yet, 37 hours after our observations, the 1st mode has evolved and increased its contribution, leading to a different lightcurve morphology between the 2 visits, which is not explained by a simple shift in phase.
We explored a method of extracting and retrieving spectra using spherical harmonics mapping at different wavelengths. This method is valuable in inferring chemical abundances and thermal profiles of potential hot and cold spots of the brightness maps. We performed spectral retrievals on 6 regions of interest but the results are not consistent. This is due to the degeneracy of 2D mapping. Observations spanning multiple rotations or Doppler tomography could decrease these degeneracies.
We also use JWST medium resolution spectroscopy for Doppler tomography of SIMP 0136. These measurement could eventually help constrain brightness maps and improve mapping fidelity. The search for a discrepancy between Doppler signal and lightcurve Fourier modes would benefit from higher spectral resolution observations.
Appendix A PCA explained variance and BIC
Appendix B Results of the regional spectra retrieval
To avoid uncorrelated noise at individual wavelengths we average the spectra in wavelength bins, then construct maps from lightcurves of binned data. We created maps from 400 light curves of 0.0045 bins using 4th order spherical harmonics. We then select 6 regions of interest and extract the spectra from 400 maps at respective longitude and latitude. The resulting spectra and a map with regions of interest are shown on Fig.10.
We then apply the same atmospheric retrieval framework as described in Section 5 to explore T-P profiles, molecular abundances and cloud levels of each region. As seen from Fig.11 the temperature profiles retrieved from regional points tend to be more isothermal in the deep atmosphere, and fail to reproduce the adiabatic slope seen in the T-P profile from the time-average integrated spectrum retrieval in Section 5. As a result, the altitudes and fractions of forsterite clouds vary significantly for each region as well. We also detect an anti-correlation, where the darkest cloud patch at lat: 0, lon: -60 has the hottest T-P profile (cyan curve in Fig.11), while the hottest point at lat: 0, lon: 15 has the lowest T-P profile (grey curve). Molecular abundances are also very different for spectrum of each region (see Figure 12).